Using the terms bonding, energy and stability, explain occurring compounds, but only 90 naturally occurring elements. why there are millions of naturally Who Gives a Darn? I Do! In completing this activity, you have explored why hydrogen atoms combine to form diatomic hydrogen, and why helium atoms do not combine to form diatomic helium. In the Star Trek universe, star ships are propelled with dilithium crystals. Dilithium is possible to prepare, however, it is only known to exist in the gas phase. To understand how this is possible, you must look at the linear combination of atomic orbitals to determine ifa stable bond can form. 1. a. How many electrons does a lithium atom have in total and in its valence shell? Only taking into account valence electrons, show how dilithium (Li2) crystals are theoretically possible b. c. What type of bond would you expect to form between two lithium atoms? Cuclant bond
media%2Fe49%2Fe498d878-ce94-4a31-8108-f5
Answer #1

a. atomic number of lithium is 3. so the total electron number is 3. the electronic configuration for Li is as follow

1S22S1 , the valence shell has only one elctron.

b. both Li atoms have following electronic configuration or atomic orbital configuration 1S22S1

but the molecular orbital configuration for Li2 is as follow

(\sigma _{1s})^{2}(\sigma^{*} _{1s})^{2}(\sigma _{2s})^{2}(\sigma^{*} _{2s})^{0}

where \sigma is bonding molecular orbital and \sigma^{*} is antibonding orbital

the standard formula to calculate the number of bonds between two atom is

bond order= ( number of bonding orbital electrons - number of antibonding electrons) / 2

now, taking consideration of only valence electron

bond order = (2 - 0) / 2 = 1......[ here, considaring only (\sigma _{2s})^{2}(\sigma^{*} _{2s})^{0} ]

it can be concluded the there is only one between two lithim atom and it is the theoretical prove of existance of dilithim crystal.

c) this is the covalent bond because this bond is formed by shareing the electrons between two lithim atom.

2. standard molecular orbital configuration for He2

(\sigma _{1s})^{2}(\sigma^{*} _{1s})^{2}

for, He2+   (\sigma _{1s})^{2}(\sigma^{*} _{1s})^{1}

so, the bond order is = (2-1) /2 = 0.5 = positive value

the value of bond order is fraction but positive, this ion can be formed in nature but it will be very unstable due to not having proper bond. therefore, it is very rare.